OpenStudy (anonymous):

PLZ HELP......you can verify the zeros of the function y=x^2+6x-7 by using a graph and finding where the graph A. is at a minimum B. is at a max C. crosses the x-axis D. crosses the y-axis

OpenStudy (anonymous):

The zeros of a function is where it crosses the x axis

OpenStudy (ksaimouli):

it is c

OpenStudy (anonymous):

thank u

OpenStudy (ksaimouli):

:-)

OpenStudy (anonymous):

mind helping me again pl

OpenStudy (ksaimouli):

ok

OpenStudy (anonymous):

find the vertex of the parabola by completing the square x^2-6x+8=y

OpenStudy (ksaimouli):

do u need the answer or explanation

OpenStudy (ksaimouli):

◊ Group ◊ y = (x² - 6x) + 8 ◊ Complete the square ◊ y = (x² - 6x + 9) + 8 - 9 ◊ Write as a binomial squared ◊ y = (x-3)² - 1

OpenStudy (anonymous):

both plz

OpenStudy (anonymous):

the choose are A.(-3,-6) b(3,-1) C(-3,1) D(-1,3)